# Fooling the Landauer bound with a demon biased thermal bath

### v1

@@ I Introduction
To challenge the generality of the Landauer principle and explore the possibility of erasing information with a lower cost [36], some underlying hypotheses in the description of the memory need to be relaxed. Notably, the effect of non-equilibrium features has been investigated on several levels [12]. To decrease the overhead to $L_{0}$ of finite-time erasures for instance, it has been theoretically proposed [41, 42] and experimentally demonstrated [35] to use out-of-equilibrium final states via adequate protocols. The Landauer bound itself can be challenged by preparing an out-of-equilibrium initial state, as suggested in Ref. 23 and demonstrated with levitated optomechanics [7]. A different approach consists in using an out-of-equilibrium thermostat: a squeezed thermal bath could potentially decrease the bound to zero [22]. For open quantum systems, the information backflow from the system to the environment leads to a memory effect and a non-Markovian dynamics, potentially altering (reducing or increasing) the Landauer bound [5].
In this work, we extend previous studies performed with a classical underdamped micro-resonator used as a one-bit memory in a feedback-controlled virtual double-well potential [8, 9]. The underdamped regime offers new insights on a wide variety of fundamental questions on the connections between feedback and thermodynamics [49, 21, 17, 44, 1], the effect on entropy and information in the presence of a continuous feedback [45, 32, 33, 6] introducing the concept of mutual information, or the theory around the Sagawa demon [46]. When the outcome of random process is stored in an external agent (demon), systems can seemingly violate the second law of thermodynamics [19]. As in the Szilard’s engine [51], the thermodynamic paradox is solved when taking into account the restoring of the external agent to its initial state [2].
@@ II Experimental set-up
Information processing is performed by applying a time protocol on the potential $U(x,t)$. A 1-bit erasure to reset the memory to state 0 in a time $2\tau$ is sketched in Fig. 1(c) [9, 8]: starting from and returning to $U_{D}(x,X_{1})$, we apply
$U(x,t)=\begin{cases}U_{D}\big[x,X_{1}\!\left(1-t/\tau\right)\big],\quad t<\tau,\\[4.0pt] U_{S}\big[x,X_{1}\!\left(t/\tau-1\right)\big],\quad\tau<t<2\tau.\end{cases}$ (1)
To explore the quasi-static regime, we typically set $\tau=$1\text{\,}\mathrm{s}$\approx 8000\,\omega_{0}^{-1}$, much longer than the oscillator relaxation time $\tau_{\mathrm{relax}}=2Q_{f}\omega_{0}^{-1}=$2.5\text{\,}\mathrm{m}\mathrm{s}$\$. The cost of information processing is measured through the stochastic work and heat $\mathcal{W}$ and $\mathcal{Q}$ defined in Ref. 50 and used in previous publication [9].
@@ IV Model: demon thermal bath
$\mathcal{W}=\Delta U_{h}=\frac{1}{2}k\big[(h-x_{1})^{2}-(h+x_{1})^{2}\big]=-2kx_{1}h.$ (2)

### v2

@@ I Introduction
To challenge the generality of the Landauer principle and explore the possibility of erasing information with a lower cost [37], some underlying hypotheses in the description of the memory need to be relaxed. Notably, the effect of non-equilibrium features has been investigated on several levels [13]. To decrease the overhead to $L_{0}$ of finite-time erasures for instance, it has been theoretically proposed [42, 43] and experimentally demonstrated [36] to use out-of-equilibrium final states via adequate protocols. The Landauer bound itself can be challenged by preparing an out-of-equilibrium initial state, as suggested in Ref. 24 and demonstrated with levitated optomechanics [7]. A different approach consists in using an out-of-equilibrium thermostat: a squeezed thermal bath could potentially decrease the bound to zero [23]. For open quantum systems, the information backflow from the system to the environment leads to a memory effect and a non-Markovian dynamics, potentially altering (reducing or increasing) the Landauer bound [5].
In this work, we extend previous studies performed with a classical underdamped micro-resonator used as a one-bit memory in a feedback-controlled virtual double-well potential [8, 9]. The underdamped regime offers new insights on a wide variety of fundamental questions on the connections between feedback and thermodynamics [50, 22, 18, 45, 1], the effect on entropy and information in the presence of a continuous feedback [46, 33, 34, 6] introducing the concept of mutual information, or the theory around the Sagawa demon [47]. When the outcome of random process is stored in an external agent (demon), systems can seemingly violate the second law of thermodynamics [20, 12]. As in the Szilard’s engine [52], the thermodynamic paradox is solved when taking into account the restoring of the external agent to its initial state [2].
@@ II Experimental set-up
Information processing is performed by applying a time protocol on the potential $U(x,t)$. A 1-bit erasure to reset the memory to state 0 in a time $2\tau$ is sketched in Fig. 2(a) [9, 8]: starting from and returning to $U_{D}(x,X_{1})$, we apply
$U(x,t)=\begin{cases}U_{D}\big[x,X_{1}\!\left(1-t/\tau\right)\big],\quad t<\tau,\\[4.0pt] U_{S}\big[x,X_{1}\!\left(1-t/\tau\right)\big],\quad\tau<t<2\tau.\end{cases}$ (1)
To explore the quasi-static regime, we typically set $\tau=$1\text{\,}\mathrm{s}$\approx 8000\,\omega_{0}^{-1}$, much longer than the oscillator relaxation time $\tau_{\mathrm{relax}}=2Q_{f}\omega_{0}^{-1}=$2.5\text{\,}\mathrm{m}\mathrm{s}$\$. The cost of information processing is measured through the stochastic work and heat $\mathcal{W}$ and $\mathcal{Q}$ defined in Ref. 51 and used in previous publication [9].
@@ IV Model: demon thermal bath
$\mathcal{W}=\Delta U_{h}=\frac{1}{2}\big[(h-x_{1})^{2}-(h+x_{1})^{2}\big]=-2x_{1}h.$ (2)